# PlanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations

### v1

@@ B.4 Room Statistics
Metric Kitchen Bedroom Living Room
Avg. # of Objects 13.30 14.24 15.29
Avg. # of Overlaps 0.52 6.27 6.26
Avg. Object Density 1.15 0.84 0.71
Appendix C Full Breakdown by Room Type, Question Category, and Model
Table 8 reports per-model accuracy separated by room type (kitchen, living room, bedroom) and question category. This is the full unaggregated results matrix underlying the analyses presented in Section 4. The table includes all eleven evaluated models and eleven question types. Path is assessed using two complementary metrics: validity and Fréchet Distance. Together, these question types span a range of challenges, covering both reasoning-heavy and functionally grounded tasks.
The table presents accuracy for all eleven models across eleven question types, grouped by room type. Each room type–kitchen, bedroom, and living room–includes results for all question types, totaling 33 rows. Each value reflects model accuracy over a fixed set of 600 questions, corresponding to 600 layouts per room type. For every combination of question type and room type, the same set of 600 questions is used across all models to ensure fair comparison.
Table 8: Question-level accuracy by category for each model across room types. Models (left to right): Qwen3-32B, DeepSeek-R1, DeepSeek-V3, LLaMA 3.3-70B-Instruct, Gemma 2-27B-it, Phi-4, GPT-4.1, LLaMA 3.1-8B-Instruct, Gemma 2-9B-it, Phi 3.5-mini, GPT-4o-mini.
Note: High truncation rate due to token limits.
Reasoning Big Models Small Models
N 32.8B 671B 671B 70B 27.2B 14B N/A 8B 9.2B 3.8B N/A
Temp. 0.6 0.6 0.3 0.0 0.5 0.0 1.0 0.6 0.5 0.7 1.0
Type Category [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
K Distance 97.2 99.5 98.5 97.2 94.7 90.7 100 82.3 74.2 35.3 97.3
Area (counters) 85.3 99.5 95.7 81.7 46.5 79.2 98.5 30.8 37.5 9.8 80.3
Free Space 83.7 88.0 52.5 37.0 23.5 42.0 83.5 9.5 13.5 6.3 20.3
View Angle 74.0 58.2 69.2 75.2 12.7 52.8 86.3 11.5 9.2 8.2 42.2
Repositioning 91.5 96.3 69.8 48.7 14.8 22.7 85.8 3.7 6.2 11.5 40.7
Max Box 33.3 30.7 9.2 8.8 3.0 3.0 57.2 1.7 1.2 1.7 3.7
Fit/Placement 92.8 92.5 71.0 66.2 72.0 82.0 89.8 68.7 71.8 68.8 70.7
Path (Valid) 13.2 6.3 39.3 30.3 23.2 26.8 73.0 10.5 17.5 14.8 35.7
Path (Fréchet) 15.0 6.3 32.8 30.2 26.0 26.0 55.8 10.2 19.0 9.3 36.7
Missing Object 87.3 88.7 44.3 56.2 52.3 52.2 79.8 27.5 39.7 14.3 58.0
Obstruction 84.0 95.2 32.7 6.0 3.3 9.7 93.5 1.2 2.7 11.3 14.2
L Distance 98.7 99.8 99.5 98.2 96.3 98.8 99.8 87.0 81.8 58.8 98.5
Area (sitting) 96.7 99.5 84.5 98.0 83.8 88.7 99.5 32.5 41.3 12.8 85.7
Free Space 0.3 4.3 1.0 0.3 3.7 0.3 5.0 0.7 3.2 1.5 1.7
View Angle 81.5 86.0 70.0 76.8 14.3 50.3 96.0 14.8 8.3 10.5 42.2
Repositioning 80.5 93.0 45.0 33.5 12.8 19.2 71.5 6.5 5.7 6.3 29.3
Max Box 1.5 0.8 1.0 3.0 2.0 2.8 6.5 0.8 1.5 1.0 2.3
Fit/Placement 90.7 91.2 71.5 80.0 83.7 87.7 91.8 75.0 75.5 72.8 72.7
Path (Valid) 10.0 14.7 33.0 26.7 30.5 26.7 53.2 7.3 21.7 16.7 33.5
Path (Fréchet) 13.2 14.2 23.7 17.8 19.2 14.7 48.0 2.5 14.2 6.8 25.2
Missing Object 73.0 76.2 51.0 49.3 29.5 36.0 65.5 9.7 28.3 11.7 32.3
Obstruction 80.7 96.5 24.3 7.3 3.8 9.3 84.7 2.5 5.2 4.7 11.7
B Distance 98.7 99.8 99.5 98.2 96.3 98.8 99.8 87.0 81.8 58.8 98.5
Area (storage) 98.7 99.8 94.0 97.0 86.3 88.3 99.3 30.3 66.3 47.7 88.2
Free Space 1.7 5.8 1.2 0.3 2.5 1.2 2.8 1.8 1.0 1.0 1.2
View Angle 76.0 79.8 70.0 78.3 10.8 57.0 94.2 15.3 10.2 7.7 43.0
Repositioning 78.7 94.3 53.8 36.0 11.2 15.7 73.7 4.3 9.0 6.5 31.5
Max Box 0.7 1.0 2.0 1.8 2.0 2.5 7.2 1.0 0.7 1.0 2.0
Fit/Placement 86.3 86.3 65.8 66.7 66.5 73.0 82.0 66.0 63.8 64.7 66.2
Path (Valid) 15.5 20.8 49.7 40.7 38.5 35.8 67.8 11.3 34.5 20.5 48.8
Path (Fréchet) 15.3 21.5 30.3 30.3 27.0 19.0 49.5 3.7 19.7 8.5 36.3
Missing Object 64.3 65.2 39.8 40.3 33.0 25.2 61.0 14.8 19.7 8.3 34.5
Obstruction 87.2 95.7 32.0 4.3 1.3 5.8 89.5 2.2 5.3 9.3 12.2
Appendix D Additional Ablations and Analysis
@@
To further verify the robustness and generalizability of our evaluation setup, we conducted several supplementary analyses beyond the primary results in Table 8.
Mean and Variability Across Room Types
Figure 6 shows radar plots summarizing model performance across question types, averaged by room type. Kitchens, living rooms, and bedrooms demonstrate similar high-level trends: models are generally most accurate on questions involving spatial metrics such as Distance, Area, and Fit/Placement, while struggling more with abstract or multi-object tasks such as Max Box and Path (Valid).
Figure 6 illustrates the standard deviation across models per question category. Notably, Repositioning, View Angle, Area, and Obstruction show the highest variability across models, particularly in bedrooms and living rooms. Free Space, in contrast, exhibits high variance only in kitchens, highlighting that performance inconsistency is not uniform across room types.
Figures 8 and 8 provide complementary radar plots organized by model, highlighting differences in room-specific performance. Kitchens consistently exhibit slightly higher mean accuracy across models, particularly for larger models like GPT-4.1 and DeepSeek-R1. Variance across models is relatively uniform, though larger models tend to show reduced variance overall.
Figure 5: Mean accuracy by question type across room categories. Each axis corresponds to a specific spatial reasoning skill (e.g., distance, placement, path validity).
Figure 6: Standard deviation in model accuracy across question types and room categories. Higher values indicate greater variability across models within a room type.
Figure 7: Mean accuracy across models by room category. Shaded regions reflect performance in bedrooms, kitchens, and living rooms, highlighting which models generalize best across spatial tasks.
Figure 8: Standard deviation of model performance across room types. Captures how consistent each model is across different spatial environments. Narrower shapes indicate more uniform behavior.
@@
To assess model robustness to prompt variation, we conducted an ablation in which each question was regenerated using the same template but with alternate object references. For example, a question originally referring to a "sofa" might instead use a "bookshelf" in the regenerated version. This allowed us to evaluate whether model performance remains stable under changes in object content while preserving linguistic structure.
As shown in Table 9, performance generally remained consistent, with only minor differences across models. Larger models such as GPT-4.1 were more stable, while smaller models like GPT-4o-mini showed slightly higher sensitivity. These results suggest that the evaluation framework is robust to this type of variation.
@@
Table 9: Accuracy under prompt variation. Each cell shows performance on the original prompt followed by a regenerated version with alternate object references.
Model Repositioning View Angle Obstruction
GPT-4.1 77.0 → 75.3 92.2 → 91.4 75.5 → 74.9
GPT-4o-mini 33.8 → 32.9 42.4 → 46.7 12.2 → 14.1
DeepSeek-R1 83.6 → 95.7 74.7 → 72.1 95.7 → 95.6
Layout Perturbation Ablation
To test model sensitivity to small physical changes, we introduced random spatial jitter of up to ±0.1 m to all object coordinates during evaluation. This simulated minor but plausible noise in the layout. After jittering, we recomputed the correct answers for all QA pairs based on the updated room configuration. The questions remained unchanged, but all evaluations used the perturbed layouts.
The jitter was constrained to ensure that objects did not overlap, remained within room boundaries, and did not block doors. Jittering can affect spatial relationships such as clearances, alignments, and angles. For example, a path that was previously clear might become blocked, or an object may no longer fit within a gap.
As shown in Table 10, model accuracy showed slight fluctuations in response to the spatial changes. Overall, model performance remained robust, indicating resilience to small, physically realistic layout variations.
Table 10: Accuracy before and after layout jitter. Values shown for three spatially sensitive questions.
Model Repositioning View Angle Obstruction
GPT-4.1 77.0 → 72.5 92.2 → 87.3 75.5 → 80.2
GPT-4o-mini 33.8 → 28.4 42.4 → 42.5 12.2 → 11.8
DeepSeek-R1 83.6 → 89.8 74.7 → 70.3 95.7 → 97.1
Appendix E Spatial Reasoning Question Suite
To extend our taxonomy of question types introduced in Table 2, Table 11 details the specific evaluation procedures used to compute ground truth answers for each category. In addition to categorizing the task, the table shows how answers are derived and what types of spatial reasoning each task demands. This breakdown clarifies not only what is being tested, but how it is measured, supporting both reproducibility and deeper analysis of model performance in later sections.
Table 11: Categories of spatial reasoning questions, answer-checking procedures, and diagnostic insights.
Category Sample Question Answer Checking Diagnostic Insight
Distance What is the Euclidean distance between the center of the door and the fridge? Compute centroids $(x,y)$ of both boxes; calculate Euclidean distance between them; compare with model output allowing a tolerance of $2\%$ of the ground-truth distance. Tests numeric computation, coordinate extraction, and arithmetic precision.
Area What is the area of the kitchen counter? Multiply the width $\times$ height of each cabinet; sum all resulting areas; compare the total against the model output with a 2% tolerance. Probes multiplicative reasoning, aggregation, and unit consistency.
Free Space What is the total unoccupied floor area in the bedroom? Merge overlapping object bounding boxes; sum the resulting areas; subtract from total room area; compare the unoccupied area with model output using a 2% tolerance. Measures the ability to aggregate and subtract multiple regions.
View Angle What is the dot product between the facing vector of the sofa and the vector to the TV? Compute normalized facing vector of reference; get normalized vector to target center; compare dot product with model output using 2% tolerance. Evaluates directional reasoning, vector normalization, and spatial alignment understanding.
Repositioning How far can the sofa move right before hitting a wall or another object? Simulate axis-aligned movement; measure distance to furthest valid non-overlapping position; compare with model output using 2% tolerance. Tests spatial simulation, collision detection, and directional constraint reasoning.
Max Box What is the largest axis-aligned box you can place in the living room? Use brute-force to solve largest empty rectangle in free space; compare resulting dimensions with 2% tolerance. Probes continuous-space optimization and spatial coverage estimation.
Fit/Placement Can a $2\,\mathrm{m}\times 3\,\mathrm{m}$ table fit in the room without overlap? Attempt to place a rectangular object in free space; ensure no intersection with existing bounding boxes. Compare boolean result with model prediction. Tests collision detection, spatial constraint enforcement, and free-space reasoning.
Shortest Path What is the shortest walkable path from the fridge to the exit with $15\,\mathrm{cm}$ clearance? Offset free space by $15\,\mathrm{cm}$ for clearance; use A* algorithm on navigable grid to compute path; compute Fréchet distance (threshold $0.6\,\mathrm{m}$) to model path; check its validity. Tests path planning, clearance-aware modeling, and heuristic graph search.
Missing Object Where should the missing armchair be placed in this layout? Check if predicted bounding box fits in free space without overlaps and satisfies clearance/accessibility constraints. Evaluates spatial reasoning, free-space utilization, and functionally appropriate object placement.
Obstruction Which objects obstruct the line of sight between the bed and the window? Compute vector from target to reference center; check which bounding boxes intersect the segment (excluding endpoints). Tests geometric reasoning, line–rectangle intersection, and occlusion detection.
Appendix F Failure Cases

### v2

@@ Appendix D Room Statistics
Metric Kitchen Bedroom Living Room HSSD
Avg. Area (m2) 12.00 17.76 20.75 17.95
Avg. # of Objects 10.35 10.76 11.69 12.20
Avg. # of Overlaps 0.52 1.82 1.52 4.39
Avg. Object Density 0.95 0.66 0.57 0.83
Avg. Vertices per Layout 41.39 43.03 46.77 152.29
To further illustrate these statistics, Figure 4 shows the distribution of object counts across all layouts. The histograms confirm the averages reported in Table 5: kitchens are concentrated at lower counts, typically around 10 objects; bedrooms and living rooms exhibit broader distributions with slightly higher counts; and HSSD layouts overlap in range but extend to higher counts in the tail. Overall, the object distributions remain comparable across sources, with no extreme outliers.
Figure 4: Distribution of object counts across kitchens, bedrooms, living rooms, and HSSD layouts.
Appendix E Full Breakdown by Room Type, Question Type, and Model
E.1 Detailed Accuracy by full dataset
Tables 6 and 7 report the full per-model accuracy matrix for the base and reasoning model groups, respectively. Each table covers nine question categories across four room types, resulting in a total of 36 rows. The path task is assessed using two complementary metrics: validity and Fréchet Distance. Together, these question types span a range of challenges, covering both reasoning-heavy and functionally grounded tasks. The base group includes eight general-purpose models, while the reasoning group includes seven models with explicit reasoning capabilities. For every model, question type accuracy is computed over a fixed set of 600 synthetic layouts (Kitchens, Living Rooms, Bedrooms) and 200 HSSD layouts, ensuring that results are directly comparable across models.
E.2 Token-Limit Analysis and Valid-Only Accuracy
In addition to overall accuracy, we analyze two complementary aspects of model performance.
First, Tables 10 and 12 quantify the fraction of responses that were terminated due to the TOKEN LIMIT stop reason. This failure mode is particularly relevant for reasoning-oriented models, which often generate longer chain-of-thought outputs.
Second, Tables 11 and 13 report accuracy over completed answers only, excluding truncated outputs (e.g., token-limit terminations). This metric isolates models’ reasoning performance on successfully produced, well-formed responses.
Together, these analyses complement the full accuracy tables by disentangling reasoning failures from generation truncation and formatting issues.
Table 6: Question-level accuracy on full dataset by room type for standard models.
Question Room [Uncaptioned image] claude sonnet-4 [Uncaptioned image] gpt-4.1 [Uncaptioned image] Kimi-K2 Instruct [Uncaptioned image] Qwen3 Coder-480B [Uncaptioned image] Qwen3 235B [Uncaptioned image] gpt-4.1 mini [Uncaptioned image] Qwen3 30B [Uncaptioned image] Devstral Small
Pair distance K 99.8 96.5 95.7 96.8 99.2 90.7 89.5 58.8
LR 99.5 95.0 94.2 96.8 99.7 88.5 88.8 62.3
B 99.7 96.3 93.3 96.5 99.5 87.8 85.2 60.0
HSSD 88.0 56.0 75.5 66.0 67.0 37.0 44.5 56.0
Placement K 87.8 78.0 82.2 80.3 90.2 86.5 85.5 74.2
LR 80.5 69.0 73.8 83.2 89.2 75.2 85.7 71.8
B 68.8 59.8 67.5 70.8 76.7 68.7 77.0 56.3
HSSD 72.0 64.5 73.0 70.0 82.0 76.5 71.5 54.5
Reposi- tioning K 73.8 63.8 48.7 45.3 83.5 66.3 73.7 14.8
LR 79.3 60.3 56.7 64.5 91.3 76.8 72.2 25.0
B 71.0 55.5 48.3 59.5 79.0 72.5 70.0 21.2
HSSD 42.0 47.0 28.0 34.0 39.0 40.5 33.0 10.0
Free space K 97.8 93.2 83.0 84.2 95.0 95.2 83.8 66.2
LR 0.2 14.2 2.8 1.8 3.5 1.2 0.5 2.7
B 2.7 31.2 1.0 1.3 8.8 0.8 1.0 0.7
HSSD 35.0 16.0 24.0 22.0 17.0 15.5 7.5 6.5
Visibility K 87.7 63.2 54.5 67.0 98.3 90.2 91.5 18.5
LR 81.7 52.7 43.3 52.2 98.3 86.8 88.7 10.0
B 74.8 57.5 41.0 54.0 96.7 86.3 86.3 10.8
HSSD 46.5 20.0 22.5 26.5 70.5 52.0 45.5 9.0
View angle K 92.0 95.3 69.8 78.8 97.0 95.8 74.7 49.7
LR 87.7 93.2 59.7 75.3 94.0 93.0 72.2 40.5
B 88.0 90.2 60.3 72.8 95.0 91.2 76.8 35.8
HSSD 67.5 55.0 28.5 46.5 50.5 46.0 34.5 30.0
Max box K 47.2 31.8 32.0 26.8 65.5 27.5 26.5 4.5
LR 7.8 7.0 5.0 5.7 22.3  4.5 8.8 0.5
B 5.8 6.8 7.3 5.0 29.2 4.8 7.0 1.7
HSSD 5.0 7.5 2.5 2.0 11.5 4.5 3.0 1.0
Shortest path (valid) K 59.2 61.7 52.7 39.7 45.2 55.3 21.8 34.2
LR 53.0 51.3 42.8 37.3 44.8 47.2 20.3 30.3
B 48.5 52.2 40.7 34.7 44.2 45.8 18.5 33.5
HSSD 28.5 25.0 18.5 18.0 26.0 15.0 5.5 10.5
Shortest path (Fréchet) K 45.3 56.8 51.3 28.2 44.2 39.5 17.8 36.2
LR 24.7 42.5 27.3 12.3 34.7 26.5 9.2 15.5
B 23.0 42.2 27.7 14.2 38.0 30.5 8.2 18.3
HSSD 15.5 22.5 18.0 8.0 20.0 12.5 2.5 11.5
Table 7: Question-level accuracy on full dataset by room type for reasoning models.
Question Room [Uncaptioned image] gpt-5 [Uncaptioned image] gpt-oss 120b [Uncaptioned image] DeepSeek R1-0528 [Uncaptioned image] Gemini Flash 2.5 [Uncaptioned image] gpt-5 mini-2025 [Uncaptioned image] gpt-oss 20b [Uncaptioned image] Qwen3 30B Think.
Pair distance K 99.8 99.3 98.0 96.3 100.0 94.2 97.7
LR 98.8 99.3 99.0 96.0 99.7 93.5 97.5
B 98.3 99.5 96.8 95.5 99.7 93.8 95.3
HSSD 69.0 78.5 25.5 12.5 32.5 40.5 18.0
Placement K 84.7 92.0 89.0 59.7 90.8 85.7 68.2
LR 75.5 89.0 82.2 53.3 86.3 78.5 46.5
B 61.2 83.5 72.0 35.8 81.2 62.0 34.5
HSSD 70.0 85.0 79.0 16.0 75.5 74.5 28.5
Reposi- tioning K 83.0 85.5 79.2 90.5 84.5 70.8 61.5
LR 85.5 89.8 86.2 91.2 92.8 87.8 77.0
B 77.8 83.3 83.3 85.2 84.3 78.0 69.2
HSSD 49.5 60.5 47.5 18.5 53.5 40.0 27.0
Free Space K 82.5 99.0 93.0 93.3 99.5 94.8 97.0
LR 47.0 83.3 18.3 17.7 78.5 53.2 0.0
B 50.5 87.5 34.8 33.3 82.2 74.0 1.2
HSSD 19.5 31.0 6.5 1.0 5.0 9.0 1.0
Visibility K 94.8 94.2 71.3 26.8 98.0 91.5 78.8
LR 95.2 94.0 52.0 11.3 98.0 89.3 70.8
B 94.2 92.5 53.5 11.2 95.5 89.2 64.2
HSSD 57.0 70.0 10.0 0.5 39.0 45.5 3.0
View Angle K 96.2 98.5 73.7 92.5 88.3 93.5 98.5
LR 93.3 97.3 68.2 91.5 84.5 92.0 98.3
B 95.2 98.2 75.2 93.8 86.8 91.3 98.3
HSSD 59.5 74.0 13.5 20.0 25.5 37.5 26.0
Max Box K 48.5 62.8 50.5 3.7 85.2 31.3 16.7
LR 17.3 28.2 8.0 0.0 60.3 3.8 0.8
B 13.0 30.3 11.0 0.0 61.2 5.8 0.5
HSSD 5.0 9.5 2.5 0.0 17.0 0.5 0.0
Shortest path (valid) K 64.7 64.2 12.3 3.2 52.3 43.0 14.2
LR 21.2 66.0 12.5 1.5 53.7 37.7 16.7
B 58.2 57.8 7.8 1.0 52.0 30.0 14.3
HSSD 28.5 33.5 8.5 0.0 16.5 13.5 1.0
Shortest path (Fréchet) K 56.3 55.5 11.5 3.2 50.5 42.2 14.0
LR 12.3 40.5 9.3 1.3 47.5 28.5 16.3
B 39.3 44.8 6.0 0.8 47.7 24.2 14.3
HSSD 22.5 26.5 2.5 0.0 12.0 14.0 1.0
Appendix F Additional Ablations and Analysis
@@
To further verify the robustness and generalizability of our evaluation setup, we conducted supplementary analyses beyond the main accuracy tables.
General models.
Figure 5 shows mean accuracy (top left) and variability (top right) across general models, as well as accuracy (bottom left) and variability (bottom right) by question type. Kitchens are consistently easier, while HSSD is the most challenging. Among models, Devstral Small performs noticeably worse, whereas Qwen3-30B achieves a level comparable to that of larger models. Across question types, Pair Distance and View Angle from the Metric category yield the highest accuracy, while more complex tasks such as Max Box and Shortest Path show lower scores and higher variance across rooms.
Reasoning models.
Figure 6 presents analogous plots for reasoning models. GPT-family models show stronger results overall. In contrast, DeepSeek-R1 and Gemini Flash 2.5 struggle with token limits, as these models tend to produce very long outputs according to Table 12. By question type, Repositioning and Placement are handled reliably, whereas Max Box and Shortest Path remain the most difficult as well, with high variance across rooms.
Figure 5: Radar plots for general models, showing mean and standard deviation of accuracy across (top) models and (bottom) question type.
Figure 6: Radar plots for reasoning models, showing mean and standard deviation of accuracy across (top) models and (bottom) question type.
@@
To assess model robustness to prompt variation, we conducted an ablation in which each question was regenerated using the same template but with alternate object references. For example, a question originally referring to a ”sofa” might instead use a ”bookshelf” in the regenerated version. This allowed us to evaluate whether model performance remains stable under changes in object content while preserving linguistic structure.
As shown in Table 9, accuracy under prompt regeneration is broadly stable; larger models change little, and smaller ones fluctuate modestly. We observe somewhat higher sensitivity for Repositioning: paraphrases can implicitly select different target objects or motion directions, occasionally introducing additional complexity or non-movable cases. Overall, the evaluation appears robust to prompt-level variations.
@@
Table 8: Accuracy under prompt variation. Each cell displays performance on the original prompt, followed by a regenerated version with alternative object references.
Model Repositioning View Angle Visibility
GPT-OSS-120B 60.5 → 59.0 74.0 → 80.0 70.0 → 77.0
GPT-OSS-20B 40.0 → 50.0 37.5 → 35.5 45.5 → 50.0
Qwen3-235B-A22B 39.0 → 49.0 50.5 → 48.0 70.5 → 74.0
Semantic Ablation
To evaluate the model’s reliance on semantic object labels rather than purely geometric reasoning, we conduct a semantic ablation experiment. In this setting, object identifiers in the scene are permuted (e.g., swapping bed and chair labels) while keeping all geometry unchanged. The regenerated prompts thus refer to the same physical configuration but with altered object semantics, for instance, a question originally phrased as “move the chair left” becomes “move the bed left,” even though the underlying geometry is identical.
As summarized in Table 9, tasks that rely primarily on pure metric computation or topological cues, such as View Angle and Visibility, show minimal changes in accuracy, with only small fluctuations due to prompt phrasing. However, performance on the action-based Repositioning task drops sharply under semantic perturbation, indicating that the model partially grounds its reasoning in object semantics rather than spatial configuration alone. This suggests that large language models may entangle linguistic priors with geometric inference when tasks involve physical movement or interaction.
Table 9: Accuracy under prompt variation. Each cell displays performance on the original prompt, followed by a regenerated version with alternative object references.
Model Repositioning View Angle Visibility
GPT-OSS-120B 60.5 → 40.0 74.0 → 73.0 70.0 → 77.0
GPT-OSS-20B 40.0 → 28.0 37.5 → 35.5 45.5 → 44.5
Qwen3-235B-A22B 39.0 → 37.0 50.5 → 43.0 70.5 → 74.0
Appendix G Case Studies by Question Type